OpenStudy (anonymous):

My answer is b 12x2 + 50xy + 8y2 a. Simplify (2x2 + y2)(x - 2y) = ? b. Simplify (2x + y)(x2 -3xy + 2y 2 ) = ? d. Simplify (6x + y)(2x + 8y )= ?

OpenStudy (anonymous):

no c?

OpenStudy (anonymous):

No I used it it's wrong

OpenStudy (anonymous):

show me :)

OpenStudy (anonymous):

narrowed the choices

OpenStudy (anonymous):

it could not be Simplify 4x(3x2+2x+1) = ? it would make no sense

OpenStudy (anonymous):

oh yes :)

OpenStudy (anonymous):

it is not b :P

OpenStudy (anonymous):

Look at d :)

OpenStudy (anonymous):

true...true....OMG IT IS D

OpenStudy (anonymous):

correction has the same (a+b)*(c+d) patter a*c=12x^2 the first number in your expression.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

not only that but b*d=8y2 :D